How long will it take Jafar and Kadir working together to cultivate a filed which Jafar can cultivate alone in three hours and Kadiraline in six hours?
Found 2 solutions by Jynald-, Alan3354:
Answer by Jynald-(6)


1/6 + 1/3 = 1/x
1/6+2/6 = 1/x
3/6 = 1/x
1/2 = 1/x
x = 2 hours

Answer by Alan3354(69443)

J does 1/3 per hour
K does 1/6 per hour
etc
============
There's a shortcut:
3*6/(3+6) = 2 hours